### v2

# Learning Manipulation Skills through Robot Chain-of-Thought with Sparse Failure Guidance

Abstract
Defining reward functions for skill learning has been a long-standing challenge in robotics. Recently, vision-language models (VLMs) have shown promise in defining reward signals for teaching robots manipulation skills. However, existing works often provide reward guidance that is too coarse, leading to inefficient learning processes. In this paper, we address this issue by implementing more fine-grained reward guidance. We decompose tasks into simpler sub-tasks, using this decomposition to offer more informative reward guidance with VLMs. We also propose a VLM-based self imitation learning process to speed up learning. Empirical evidence demonstrates that our algorithm consistently outperforms baselines such as CLIP, LIV, and RoboCLIP. Specifically, our algorithm achieves a 5.4×5.4$\times$5.4 × higher average success rate compared to the best baseline, RoboCLIP, across a series of manipulation tasks.
Keywords: Manipulation Skills, Vision Language Models, Chain-of-Thought
1 Introduction
Defining reward functions for learning robotic skills has been a long standing challenge in robotics [1]. Traditional methods typically define the rewards through human knowledge and require significant engineering efforts. Some methods also use expert demonstrations to define rewards [1]. However, this procedure requires expert supervision and can be infeasible in complex scenarios [2] . Recently, the success of vision language models (VLM) has inspired researchers to leverage its power to define rewards for skill learning. For example, RoboCLIP [2] captures the temporal information from robot movements with a video backbone S3D [3].
In spite of the preliminary success of VLM-based reward modeling, a downside of these existing works is that their language guidance is too coarse for learning. Consider the task of opening the door, existing works only use the language embedding of "opening the door" in reward definition to guide the learning process. Such guidance can be too sparse for learning: when the robot is not able to approach the door at the beginning of learning process, the rewarder will always provide a near-zero feedback since the door is not likely to be opened. In contrast, in human world we typically use more fine-grained guidance such as "move towards the door", "move the hand towards the handle" and "grab the handle" during the teaching process. Intuitively, using such decomposition can make the guidance more precise, and will lead to richer and denser feedback.
Based on this intuition, in this paper we present a novel framework named Rewarder from Robot Chain-of-Thought with Sparse Failure Guidance (RoboCoT) to address this problem. Our main insight is to decompose the task description into a series of actionable steps and then leverage VLM to define a fine-grained reward to guide skill learning, which can make the learning process more sample efficient and yield higher success rates. We also propose a VLM-based self-imitation procedure to further speed up the learning process. In the experiments, we demonstrate that our approach performs well across a broad spectrum of tasks, achieving a 5.4×5.4$\times$5.4 × performance improvement over the best existing baseline. Ablation studies show that each component of our algorithm contributes to the overall performance.
Figure 1: First, the robot receives the language instruction, e.g., "open the door." Subsequently, our algorithm interprets this instruction using a robot chain-of-thought processing, breaking it down into several detailed prompts. The robot also considers some failure guidance. Then, the robot initiates reinforcement learning with our designed reward signal to learn the skills. Successful experiences are recorded by foundation models to reinforce effective behaviors (i.e., VLM-based self-imitation).
2 Related Work
2.1 Learning Manipulation Skills for Robotics
Acquiring general manipulation abilities can be facilitated through learning from expert demonstrations, using behavioral cloning (BC) [4, 5, 6] or inverse reinforcement learning (IRL) [7, 8]. BC directly maps the state space to the action space to mimic the expert’s actions. However, BC faces several limitations, including compounding errors [9], the copycat problem [10], and out-of-distribution sample shifts [11]. In contrast, IRL derives both the reward function and the policy from expert demonstrations. However, IRL struggles with sample efficiency, particularly in real-world applications [12, 13, 14].
Mastering manipulation skills with language inputs directly presents several challenges, such as a lack of reward signals and inefficiency in understanding vision-language contexts. Recent studies have explored integrating VLMs to enhance the understanding of language instructions [15]. Our algorithm is designed by effectively leveraging VLMs as a source of rewards.
2.2 Foundation Models for Robotics
VLMs equip robot agents with compact pre-trained representations. R3M [16] co-trains video and language instructions in a self-supervised manner [17], enhancing the robot’s ability to mimic expert actions. PaLM-E [18] integrates visual observations with large language models, facilitating low-level robot control that includes reasoning about the control process. RT-1 [19] uses transformers to better mimic expert behaviors. RoboFlamingo [20] uses multi-step observation inputs to capture a comprehensive view of observational data. RT-2 [21] innovates by conceptualizing robot actions as a form of language. Furthermore, RT-H [22] introduces an action hierarchy through language motions, leveraging the chain-of-thought reasoning capabilities of large language models.
VLMs significantly enhance robot manipulation capabilities by offering valuable reward signals. CLIP [23], for instance, aligns image representations with language embeddings. Consequently, [24] employs CLIP to measure the reward distance between a robot’s current observation and its linguistic goal. LIV [25] innovatively combines VIP [26] and CLIP to more effectively learn both visual and linguistic goal representations. RoboCLIP [2] adopts the S3D [3] representation model to assess the reward for the final observation of a trajectory. Furthermore, RL-VLM-F [27] directly utilizes GPT-4V to label preferences for agent rollout data.
However, current models still struggle to master a broad spectrum of robot tasks, often succeeding only in simple tasks such as closing a drawer. VLMs like CLIP [23] and LIV [25] lack an understanding of the manipulation process. On the other hand, RoboCLIP [2] lacks spatial reasoning capabilities and task completion awareness. RL-VLM-F [27] excels in spatial reasoning for manipulation tasks but is deficient in process understanding capabilities. In contrast, our algorithm not only comprehends detailed language instructions from the robot chain-of-thought but also considers failure scenarios. Additionally, it leverages VLM’s task completion reasoning ability through self imitation process, offering a more comprehensive approach to master robot tasks.
3 Background
We consider scenarios where the robotic system receives specific verbal commands, such as "open the door." To model this learning process, we employ a Markov Decision Process (MDP) defined by $(S,A,r,P,\gamma)$, where $S$ denotes the state space, $A$ the action space, $r$ the reward model, $P$ the transition model, and $\gamma$ the discount factor.
Chain-of-Thought (CoT) The CoT [28] approach is a reasoning method designed to enhance the problem-solving capabilities of large language models. This method involves decomposing a problem into a series of intermediate steps or thoughts that collectively guide towards a solution, mirroring the human approach to complex problem solving. By explicitly detailing these intermediate steps, large language models such as GPT-4V are better equipped to navigate and resolve complex challenges that require advanced understanding and logical deduction.
VideoCLIP VideoCLIP [29] is a multi-modal model that connects visual and textual information, specifically designed to understand videos by associating them with relevant text descriptions. The model is trained on a diverse dataset of videos and associated texts, allowing it to develop a nuanced understanding of the relationship between video segments and language. In particular, VideoCLIP minimizes the sum of two multi-modal contrastive losses:
$\mathcal{L}=-\sum_{(v,t)\in\mathcal{B}}(\log\text{NCE}(z_{v},z_{t})+\log\text{NCE}(z_{t},z_{v})),$ (1)
where $\mathcal{B}$ is the batch containing sampled video-text pairs, and NCE$(z_{v},z_{t})$ and NCE$(z_{t},z_{v})$ correspond to the contrastive loss on video-to-text similarity and vice versa. Specifically, the video-to-text contrastive loss is given by:
$\text{NCE}(z_{v},z_{t})=\frac{\exp(z_{v}\cdot z_{t}^{+}/\tau)}{\sum_{z\in{\{z_{t}^{+},z_{t}^{-}\}}}\exp(z_{v}\cdot z/\tau)},$ (2)
where $\tau$ is a temperature hyper-parameter, $z_{t}^{+}$ and $z_{t}^{-}$ are positive and negative text embeddings respectively, and $z_{v}$ is a video segment embedding.
4 Method
Figure 2: First, text embeddings are generated using robot chain-of-thought (positive prompts) and human failure guidance (negative prompts). Next, a moving window slides across the agent’s rollout trajectory, calculating the NCE reward based on both video and text embeddings. Finally, VLMs evaluate the final observation. If this observation is deemed successful, this trajectory is recorded in the success replay buffer for self imitation.
In section 4.1, we introduce the robot chain-of-thought interpretation. In section 4.2, we explain how to define the rewards for policy learning with VLMs. Finally, in section 4.3, we introduce VLM-based self imitation procedure via success experience regularization.
4.1 Robot Chain-of-Thought
Manipulation skills consist of a series of actionable control steps. For instance, opening a door includes sub-tasks such as approaching the handle, grasping it, pulling it back, and opening the door. Decomposing tasks simplifies the manipulation process, affording robots a greater level of understanding and control.
We use GPT-4V to interpret task instructions such as "open the door." With GPT-4V, this instruction is broken down into a series of sub-goals $l_{i}^{+}$ through a CoT process, such as "the robot hand approaches the door’s handle," "the robot hand grasps the door’s handle," "the robot hand pulls the door’s handle back," and "the robot hand opens the door." Consequently, we leverage the fine-grained task descriptions from GPT-4V interpretation to enhance the agent’s understanding of the tasks.
4.2 VLMs as Rewarder
A central challenge we address is defining a reward function that effectively reinforces desired behaviors. In this work, we argue that VLMs can provide the agent with a much more fine-grained reward guidance. Consequently, we propose a novel reward design mechanism based on robot CoT. Specifically, we explore the usage of VideoCLIP [29] as a VLM. VideoCLIP excels in capturing the temporal dynamics of video segments and interpreting associated language prompts, aiming to align robot movements with actions described in the task descriptions.
The reward signal for learning manipulation skills is defined as follows:
$\text{reward}=\frac{\max_{z_{l}\in\{z_{l}^{+}\}}\exp(z_{v}\cdot z_{l}/\tau)}{\max_{z_{l}\in\{z_{l}^{+}\}}\exp(z_{v}\cdot z_{l}/\tau)+\sum_{z^{\prime}_{l}\in{\{z_{l}^{-}\}}}\exp(z_{v}\cdot z^{\prime}_{l}/\tau)},$ (3)
where $z_{v}$ is the video segment embedding obtained using a moving window sliding on the agent rollout trajectory, $z_{l}^{+}$ is a set of detailed task description embeddings from robot CoT interpretation $l^{+}$, and $z_{l}^{-}$ is a set of potential failure case description embeddings guided by humans. Here, offline human failure guidance helps the agent consider potential failure cases, improving the policy for robustly learning manipulation skills. To enable the agent to recognize and adjust to potential setbacks, we provide sparse human failure guidance before training with 2 to 4 prompts. The training of VideoCLIP, employing a methodology akin to the InfoNCE loss, aims to distinguish between video segments and corresponding positive or negative texts. We argue that a similar NCE-based reward formulation can leverage the VideoCLIP model rather than the reward definition in RoboCLIP [2]. Moreover, Equation 3 ensures that robot behaviors aligning with the detailed task descriptions and avoiding failure scenarios yield higher rewards to the agent. Otherwise, the reward signal would penalize undesired behaviors.
4.3 VLM-based Self Imitation
Our reward design mechanism motivates the agent to explore the necessary manipulation skills. Empirically, we find that the VideoCLIP rewards don’t reflect the completion of the manipulation tasks well. They focus mostly on whether the robot’s movements align with detailed language goals. To better reinforce task completion and speed up the learning, we introduce VLM-based self imitation procedure.
We employ GPT-4V to assess the success of the agent’s final observation in its rollout trajectory. If it deemed successful, this trajectory is archived into a success replay buffer and we correct the reward function as: $r(s_{T})=r(s_{T})+100$. Here, $s_{T}$ is the successful final observation of the rollout trajectory. Subsequently, we introduce a regularization mechanism aimed at refining the agent’s policy to reinforce successful behavior. This motivates the agent to reinforce the success experience. The regularization loss is integrated with the policy loss, formulated as follows:
=RL $\mathcal{L}=\text{RL Loss}+\lambda\sum_{(s,a)\in D_{E}}-\log(\pi(a|s)),$ (4)
where $D_{E}$ is the success experience replay buffer and $\lambda$ is the hyper-parameter. By incorporating success experiences into the policy loss calculation, the agent is encouraged to repeat desired behaviors.
5 Experiment
In this section, we assess the effectiveness of our proposed algorithm RoboCoT. The experiments are designed to address two questions: (1) How does our algorithm perform compared to other state-of-the-art algorithms? The results are presented in section 5.1. (2) What are the contributing factors in our algorithm? The ablation studies are presented in section 5.2.
Environments To evaluate the effectiveness of our proposed algorithm RoboCoT, we conduct experiments on a wide array of manipulation tasks using MetaWorld-v2 as the test robotic environment. We select 10 tasks from simple to complex, including drawer, window, faucet, button press, door, and soccer, as described by [30]. Among these tasks, opening objects such as drawers, windows, faucets, and doors is significantly more challenging than closing them.
Setup To ensure a fair evaluation across all experiments, we use DrQ-v2 [31] as the reinforcement learning algorithm. Detailed implementation specifics can be found in the references: CLIP [24], LIV [25], and RoboCLIP [2]. All experiments, including those involving our proposed algorithm, receive the same language instructions as inputs. Implementation details are provided in the Appendix 8.
5.1 Overall Comparison
Figure 3: Training curves for baselines and our algorithm within 3 random seeds. Each data point is evaluated with 20 sampled trajectories. The shaded area displays the range of one standard deviation. Our algorithm gets best performance compared to other baselines, achieves 5.4×5.4$\times$5.4 × improvement compared to RoboCLIP.
To demonstrate the superiority of our algorithm, we conduct a series of experiments comparing it with established baselines: CLIP [24], LIV [25], and RoboCLIP [2]. CLIP [24] employs CLIP [23] as the vision-language model to measure the distance between current observations and language goals. LIV [25] aligns goal representations with language instructions, and RoboCLIP [2] utilizes S3D [3] to measure the distance between trajectory videos and language as a form of reward. However, CLIP [23] and LIV [25] lack an fine-grained reward guidance, while RoboCLIP [2] lacks task completion awareness. Figure 3 illustrates the training curves, depicting the success rates for each algorithm. Each evaluation involves 20 trajectories, and the experiments were replicated across three different environmental random seeds to introduce variability in the position of objects.
Our findings indicate distinct performance trends across various tasks. RoboCLIP displays promising results in tasks involving drawer and window manipulation, as well as soccer, but it is most effective in the simplest tasks such as drawer closing. CLIP performs well in tasks related to faucet, drawer and window manipulation, and soccer, indicating that CLIP focuses more on task completion rather than the manipulation process. LIV shows competence primarily in tasks involving drawer closure, window opening, and soccer. Our algorithm RoboCoT consistently outperforms the baselines across all tasks, achieving success rates exceeding $75\%$ in tasks such as closing and opening drawers, windows, and doors. These results underscore the effectiveness of our approach in handling diverse and challenging scenarios.
5.2 Ablation Study
In this section, we conduct ablation studies by removing each component of our algorithm (CoT interpretation, human failure guidance, and self imitation). The results indicate that the most contributing factor to our algorithm’s performance is VLM-based self imitation procedure, followed by human failure guidance. Each component contributes to the overall performance of our algorithm. Additionally, we conduct ablation studies comparing GPT-4V success labels with ground truth success labels, demonstrating the robustness of our GPT-4V labeling. Finally, experiments with all baselines incorporating VLM-based self imitation procedure reveal that our algorithm’s superiority is not solely due to self imitation but also the reward design mechanism.
Question 1: how significant is each component of our algorithm in contributing to the results?
Figure 4: Final policy performance for ablations and our algorithm. The results show that VLM-based self imitation procedure is the most significant contributing factor in our algorithm, followed by human failure guidance.
Our algorithm incorporates several critical components: robot chain-of-thought, sparse human failure guidance, and VLM-based self imitation. To evaluate the significance of each component, we conducted a series of experiments, systematically removing each component to observe its impact on performance. A decline in performance upon the removal of any component demonstrates its essential role in our algorithm.
Figure 4 illustrates the success rates of the final policy performance of our algorithm compared to its ablated versions. The labels for these ablations are: w/oReg (without success experience regularization, i.e., self imitation), w/oHG (without sparse human failure guidance), and w/oCoT (without robot chain-of-thought task decomposition).
The results confirm the critical importance of each component in our algorithm, with VLM-based self imitation being particularly pivotal. Human failure guidance is the second most important component, motivating the agent to consider potential failure cases. Additionally, robot chain-of-thought processing helps the agent better understand the process of manipulation tasks, further contributing to the performance of our algorithm.
Question 2: what is the empirical difference between the labels assigned by GPT-4V and the ground truth labels in the context of success experience collection?
Figure 5: Ablation study for our algorithm with different success labeling methods (GPT-4V labeling or ground truth (GT) labeling). The results demonstrate that the success labels generated by GPT-4V are robust.
Figure 6: Ablation study for all baselines with success experience regularization. The results show that our algorithm performs well not only because of the VLM-based self imitation component but also due to the reward design mechanism.
Our algorithm RoboCoT incorporates GPT-4V to gather successful trajectories for VLM-based self imitation procedure. To validate the robustness of our GPT-4V labeling in collecting success experiences, we conduct experiments comparing GPT-4V labels and ground truth labels. If the performance using GPT-4V labels is similar to that using ground truth labels, it would confirm that GPT-4V is robust in collecting success experiences.
The process for GPT-4V analysis is outlined in Figure 2. Initially, we present GPT-4V with a scenario indicating a successful state (e.g., the door is open). We then provide a prompt: "Consider the following image; this is a simulation environment. In this image, the door of the box is open. We define it when the door is open, the output is 1. Otherwise, the output is 0." Subsequently, we present the final observation to GPT-4V and prompt, "What is the output of this image?" We collect the trajectory as successful when GPT-4V’s response is 1.
Figure 6 displays the training success rate curves comparing GPT-4V and ground truth (GT) labels for success experience collection. Our findings show that the GPT-4V labels achieves comparable performance to the ground truth across various tasks, demonstrating the robustness of our GPT-4V labels.
Question 3: what is the empirical difference between our algorithm and all baselines with VLM-based self imitation?
Question 1 shows that VLM-based self imitation component plays the most important role in our algorithm. To evaluate the significance of the reward design mechanism, we conduct experiments applying VLM-based self imitation procedure to all baselines. If the results show that our algorithm still outperforms other baselines with self imitation procedure, then it proves that our algorithm’s superiority is due not only to the VLM-based self imitation procedure but also to the reward design mechanism.
Figure 6 shows the ablation results comparing our algorithm with other state-of-the-art baselines that include VLM-based self imitation procedure. The results demonstrate that our algorithm outperforms other baselines on almost all tasks. This indicates that our algorithm’s superiority is due not only to the VLM-based self imitation component but also to the reward design mechanism.
6 Conclusion
In this study, we present a novel algorithm RoboCoT, which defines rewards for skill acquisition. Our algorithm RoboCoT simplifies complex tasks through a robot’s chain-of-thought process with sparse failure guidance. It yields a more fine-grained reward guidance. Through VLM-based self imitation, our algorithm strengthens its policies, thereby improving performance. Empirical results confirm the effectiveness of our approach, demonstrating a distinct advantage over established baseline models. This research sets a new standard for mastering manipulation skills in robots, and we hope it can be applied to more robotic tasks in the future.
7 Limitation
Our proposed method, RoboCoT, effectively masters a wide range of manipulation tasks through a robot chain-of-thought with sparse failure guidance. It provides a fine-grained reward guidance for the robot agent, facilitating the training of various automation tools. However, these tasks primarily focus on aligning robot movements with detailed language descriptions. RoboCoT does not yield promising results in tasks that emphasize observational preferences, such as soccer.